# time2time: Causal Intervention in Hidden States to Simulate Rare Events in Time Series Foundation Models

### v1

@@ Abstract
While transformer-based foundation models excel at forecasting routine patterns, two questions remain: do they internalize semantic concepts such as market regimes, or merely fit curves? And can their internal representations be leveraged to simulate rare, high-stakes events such as market crashes? To investigate this, we introduce activation transplantation, a causal intervention that manipulates hidden states by imposing the statistical moments of one event (e.g., a historical crash) onto another (e.g., a calm period) during the forward pass. This procedure deterministically steers forecasts: injecting crash semantics induces downturn predictions, while injecting calm semantics suppresses crashes and restores stability. Beyond binary control, we find that models encode a graded notion of event severity, with the latent vector norm directly correlating with the magnitude of systemic shocks. Validated across two architecturally distinct TSFMs, Toto (decoder only) and Chronos (encoder decoder), our results demonstrate that steerable, semantically grounded representations are a robust property of large time series transformers. Our findings provide evidence for a latent concept space that governs model predictions, shifting interpretability from post-hoc attribution to direct causal intervention, and enabling semantic “what-if” analysis for strategic stress-testing.
@@ 1 Introduction
Our results provide a striking confirmation. Injecting the statistical signature of a major crash (e.g., 2000  [^1] https://en.wikipedia.org/wiki/Dot-com_bubble, 2008  [^2] https://en.wikipedia.org/wiki/2008_financial_crisis, or 2020  [^3] https://en.wikipedia.org/wiki/2020_stock_market_crash) forces a calm period’s forecast into a sharp downturn. The converse intervention; imposing calm statistics onto a crash, suppresses the downturn forecast. Beyond this binary steering, we uncover a deeper nuance in the model’s understanding: the model quantitatively encodes the relative severity of different crashes, as measured by the norm of its latent representation. The model has learned not just what a crash is, but that the Dot-com crash of 2000 constitutes a more severe representational event than the 2008 crisis, revealing a rich, semantically grounded internal world. Building on these results, this work makes the following contributions:
@@ Appendix A Limitations and Future Work
While our work provides a foundational proof-of-concept for semantic control, it represents the first step into an exciting research landscape. Researchers can investigate whether these steerable semantic subspaces are a universal property of sequence models, extending beyond Transformers to architectures like State Space Models Rangapuram et al. (2018); Alexandrov et al. (2019) or LSTM based time series models Wen and Li (2023); Siami-Namini et al. (2019); Lim et al. (2020). Furthermore, our intervention technique opens the door to discovering a broader “semantic vocabulary” within TSFMs, probing for concepts like earnings surprises in finance or seizure onsets in EEG data. Our findings invite a new research program focused not just on what these models can predict, but on the rich, internal worlds they build to do so.

### v2

@@ Abstract
While transformer-based foundation models excel at forecasting routine patterns, two questions remain: do they internalize semantic concepts such as market regimes, or merely fit curves? And can their internal representations be leveraged to simulate rare, high-stakes events such as market crashes? To investigate this, we introduce activation transplantation, a causal intervention that manipulates hidden states by imposing the statistical moments of one event (e.g., a historical crash) onto another (e.g., a calm period) during the forward pass. This procedure deterministically steers forecasts: injecting crash semantics induces downturn predictions, while injecting calm semantics suppresses crashes and restores stability. Beyond binary control, we find that models encode a graded notion of event severity, with the latent vector norm directly correlating with the magnitude of systemic shocks. Validated across two architecturally distinct TSFMs, Toto (decoder only) and Chronos (encoder-decoder), our results demonstrate that steerable, semantically grounded representations are a robust property of large time series transformers. Our findings provide evidence for a latent concept space that governs model predictions, shifting interpretability from post-hoc attribution to direct causal intervention, and enabling semantic “what-if” analysis for strategic stress-testing.
footnotetext: Code available at https://github.com/birla-ai-labs/time2time**footnotetext: Correspondence: saurabh.deshpande-c@adityabirla.com
@@ 1 Introduction
Our results provide a striking confirmation. Injecting the statistical signature of a major crash (e.g., 2000 https://en.wikipedia.org/wiki/Dot-com_bubble, 2008 https://en.wikipedia.org/wiki/2008_financial_crisis, or 2020 §§§https://en.wikipedia.org/wiki/2020_stock_market_crash) forces a calm period’s forecast into a sharp downturn. The converse intervention; imposing calm statistics onto a crash, suppresses the downturn forecast. Beyond this binary steering, we uncover a deeper nuance in the model’s understanding: the model quantitatively encodes the relative severity of different crashes, as measured by the norm of its latent representation. The model has learned not just what a crash is, but that the Dot-com crash of 2000 constitutes a more severe representational event than the 2008 crisis, revealing a rich, semantically grounded internal world. Building on these results, this work makes the following contributions:
@@ Appendix A Limitations and Future Work
While our work provides a foundational proof-of-concept for semantic control, it represents the first step into an exciting research landscape. Researchers can investigate whether these steerable semantic subspaces are a universal property of sequence models, extending beyond Transformers to architectures like State Space Models Rangapuram et al. (2018); Alexandrov et al. (2019) or LSTM based time series models Wen and Li (2023); Siami‐Namini et al. (2019); Lim et al. (2020). Furthermore, our intervention technique opens the door to discovering a broader “semantic vocabulary” within TSFMs, probing for concepts like earnings surprises in finance or seizure onsets in EEG data. Our findings invite a new research program focused not just on what these models can predict, but on the rich, internal worlds they build to do so.